OpenStudy (anonymous):

Find the inverse: y=x^2+2x+1 please :)

OpenStudy (anonymous):

replace x by y and y by x x = y^2 + 2y + 1 x = (y+1)^2 y+1 = +/- sqrt(x) \[y=-1\pm \sqrt{x}\] replace y by \[y ^{-1} = -1\pm \sqrt{x}\]

OpenStudy (anonymous):

read it as y inverse

OpenStudy (anonymous):

ah beautiful! thank you.